Fe2O3( Molar mass= 159.7g/mol) reacts with CO (molar mass=28.0 g/mol), according to the equation Fe2O3(s) + CO (g) = CO2 (g) + Fe (s). When 116g of CO reacts with Fe2O3. How many moles of iron(Fe) will be produced?


1
Expert's answer
2021-02-24T03:38:22-0500

Fe2O3(s) + 3 CO (g) = 3 CO2 (g) + 2 Fe (s)

Given mass of CO is greater than that taken by the reaction: 116/23/3 = 1.38 times. The produced moles of Fe is: 2*1.38 = 2.76